Let X be a non-empty set of real numbers. If it has a greatest lower bound, then it must be unique


1
Expert's answer
2021-01-18T05:43:40-0500

The Trichotomy Principle for real numbers states that for "x,y\in\mathbb R" either "x<y" or "x=y" or "x>y".


Let "X" be a non-empty set of real numbers. Suppose there exists two such greatest lower bounds "a,b\in\mathbb R"  such that "a\le x"  and "b\le x" for all "x\in X". So "a" and "b" are both lower bounds for the set "X" and in particular, since "a"  and "b"  are both greatest lower bounds, "a\le b"  and  "b\le a"  by the definition of a greatest lower bound. Hence by the Trichotomy Principle, we conclude "a=b". Therefore, a greatest lower bound must be unique.